OpenStudy (erinweeks):

Find all the zeros of the equation. -4x4 - 44x2 + 3600 = 0

OpenStudy (erinweeks):

A. 5, -5, 6i, -6i B. 5, 6i C. 5, -5, 6i, 0 D. -5, -6i

OpenStudy (campbell_st):

divide every term by -4 so \[x^4 + 11x^2 - 900 = 0\] this is basically a quadratic so you can do a substitution and let \[x^2 = u\] then you have the equation \[u^2 + 11u - 900 = 0\] this may make it easier for you to solve.

OpenStudy (erinweeks):

so i sub -4 in for u and divide?

OpenStudy (campbell_st):

nope... I've done the division... you have \[x^4 + 11x^2 - 900 = 0\] find some factors of -900 that add to 11... try -25 as 1.... you can find the other...

OpenStudy (erinweeks):

im confused? would it be D?

OpenStudy (campbell_st):

nope... just do some math and you'll get the solution...

OpenStudy (erinweeks):

it makes no sense to me i dont get how there getting the solutions

OpenStudy (campbell_st):

ok... so how many -25 in - 900...?

OpenStudy (erinweeks):

how many what ??

OpenStudy (erinweeks):

36

OpenStudy (campbell_st):

ok... so if you factorise your polynomial you have \[(x^2 - 25)(x^2 + 36) = 0\] you need to solve the quadratics \[x^2 - 25 = 0\] and \[x^2 + 36 = 0\] I'll let you do that.

OpenStudy (erinweeks):

Well i dont know the answer but x^2-25 = 0 5^2 - 25 = 0 and x^2 + 36 = 0 -6^2 +36 = 0

OpenStudy (erinweeks):

Im not sure what the answer would be.. A?

OpenStudy (erinweeks):

@jim_thompson5910 please help !!

jimthompson5910 (jim_thompson5910):

if there are complex roots, then they ALWAYS come in conjugate pairs

jimthompson5910 (jim_thompson5910):

so if you had 2i as a root, then -2i would automatically be another root

jimthompson5910 (jim_thompson5910):

the only place where this happens is choice A, so you're correct

OpenStudy (erinweeks):

So it is A right? i did my work correctly to?

jimthompson5910 (jim_thompson5910):

yes it looks good

OpenStudy (erinweeks):

thank you Jim

jimthompson5910 (jim_thompson5910):

np